PRESSURE BROADENING AND SHIFT OF SOME NH_{3} LINES MEASURED BY DIODE LASER SPECTROSCOPY

\underline{KOICHI M.T. YAMADA} , M. FABIAN , National Institute for Advanced Interdisciplinary Research, Tsukuba 305, Japan ; R. SCHIEDER, and G. WINNEWISSER , I. Physikalisches Institut, Universität zu Köln, D-50937 Köln, Germany .

Using the stabilized diode laser spectrometer in Cologne we have precisely determined pressure-induced line broadening and line shift coefficients of six rovibrational transitions in the ν_{2} band of NH_{3}. Foreign gas effects due to O_{2}, N_{2}, H_{2}O, CO, He, and Ar have been investigated and the self-effects as well.

The high accuracy in the frequency position of the laser was achieved by stabilizing the diode laser on an interference maximum of a Fabry-Perot interferometer of variable optical length. The laser frequency was swept by tuning the cavity length of the étalon; the cavity length was monitored by a stabilized HeNe laser. A simultaneous recording of an NH_{3} spectrum with a reference cell at a fixed pressure has been introduced for an absolute calibration. The line positions relative to reference lines have been measured to a precision of 300 kHz. A good signal-to-noise ratio was achieved by using a Herriott-type multi-reflection cell with a total path length of 46.86 m. This allowed to use very small NH_{3} partial pressures of the order of a few hundred μbar in the measurements of foreign gas effects to minimize the influence of NH_{3} self-effects.

It turned out that contamination of the gas sample with H_{2}O imposed a serious problem upon the precise determination of the pressure effects, because of the large shift and broadening caused by H_{2}O. To remove any residual water contamination, the sample cell was connected to a cold finger submerged in a Dewar vessel filled with dry ice. The results of the present study are thus thought to be the most precise data of this kind now available.